OpenStudy (anonymous):

What is the simplest form of this expression? (x^-3)^1/3 a. x^-1 b. x^9 c. 1/x^9 d. 1/x

OpenStudy (anonymous):

A.

OpenStudy (anonymous):

-3 reduced by 1/3 is a -1

OpenStudy (solomonzelman):

\(\Large\color{black}{ \displaystyle \left(x^{\color{blue}{\rm a}}\right)^{\color{red}{\rm b}} =x^{\color{blue}{\rm a}~\times~\color{red}{\rm b}} }\)

OpenStudy (solomonzelman):

This is the property you need. In your case: * a = -3 * * b = 1/3 *